1. A firm has a bond issue with face value of $1,000, 8% coupon rate, and eight years to maturity. The bond makes coupon payments every six months, and is currently priced at $1,055.85. What is the yield to maturity on this bond?

Select one: a. 3.54% b. 6.95% c. 7.07% d. 7.49% e. 14.99%

2. What is the duration of a five-year bond with coupon rate of 8%, yield to maturity of 6%, semi-annual coupon payment, and face value of $1,000?

Select one: a. 3.80 years b. 3.95 years c. 4.20 years d. 4.25 years e. 8.51 years

3.

The yield on a 10-year bond is 6.5%. The 30-day T-bill yield is 3.5%, while the inflation rate is estimated to be 2.5%. What is the real rate of return on the bond based on the exact Fisher Effect formula?

Select one: a. 3.00% b. 3.90% c. 4.00% d. 6.50% e. 9.16%
Answer #1

Interest rate per semiannually (8%/2) 4.00%
yield to maturity per semiannually (6%/2) 3.00%
Interest paid per semiannually (1000*4%) $                  40
Last payment (40+1000 principal) $            1,040
Period (T) Cash Flow PV factor @ 3% Present value Weight (Present value * Period)
1 $          40           0.9708738           38.83                38.83
2 $          40           0.9425959           37.70                75.41
3 $          40           0.9151417           36.61              109.82
4 $          40           0.8884870           35.54              142.16
5 $          40           0.8626088           34.50              172.52
6 $          40           0.8374843           33.50              201.00
7 $          40           0.8130915           32.52              227.67
8 $          40           0.7894092           31.58              252.61
9 $          40           0.7664167           30.66              275.91
10 $    1,040           0.7440939         773.86          7,738.58
Total     1,085.30          9,234.50
Duration in semi annual period (9234.50/1085.30) 8.5087
Duration in yearly period (8.5087/2) 4.25
Correct option is D
(1+Real rate of Return) = (1+ Nominal rate) / (1+Inflation rate)
(1+Real rate of Return) = (1+ 0.065) / (1+ 0.025)
(1+Real rate of Return) = 1.065/1.025
1+Real rate of Return = 1.0390
Real rate of Return = 1.0390 - 1
Real rate of Return = 0.039
Real rate of Return = 3.90%
Real rate of Return = 3.90% 3.90%
Correct option is B